how would i solve;
(x+5)^(1/2)-x=3

Answer by stanbon(48535)

solve;
(x+5)^(1/2)-x=3
---
Rearrange:
sqrt(x+5) = x+3
---
Square both sides to get:
x+5 = x^2 + 6x + 9
x^2 + 5x +4 = 0
Factor:
(x+4)(x+1) = 0
x = -4 or x = -1
-----------------
Check the potential answers in the original equation:
Check x = -4: sqrt(-4+5) = -4+3; sqrt(1) = -1 ; false
----------
Check x = -1: sqrt(-1+5) = -1+3 ; sqrt(4) = 2 ; true
----------------
Soltuion: x = -1
========================
Cheers,
Stan H.